# Evaluating Diversity in Automatic Poetry Generation

### v1

@@ Abstract
Natural Language Generation (NLG), and more generally generative AI, are among the currently most impactful research fields. Creative NLG, such as automatic poetry generation, is a fascinating niche in this area. While most previous research has focused on forms of the Turing test when evaluating automatic poetry generation — can humans distinguish between automatic and human generated poetry — we evaluate the diversity of automatically generated poetry, by comparing distributions of generated poetry to distributions of human poetry along structural, lexical, semantic and stylistic dimensions, assessing different model types (word vs. character-level, general purpose LLMs vs. poetry-specific models) and types of fine-tuning (conditioned vs. unconditioned). We find that current automatic poetry systems are considerably underdiverse along all dimensions — they tend to memorize, do not rhyme sufficiently, are semantically too uniform and even do not match the length distribution of human poetry. Among all models explored, character-level style-conditioned models perform slightly better. Our identified limitations may serve as the basis for more genuinely creative future poetry generation models.
@@
Anonymous ACL submission
@@ 1 Introduction
A key aspect of creative language generation is the ability to create new, original and interesting text, cf. colton2012computational; gatt2018survey; yi2020mixpoet; Elgammal2017CANCA. To date, extremely little attention has been given to the evaluation of originality and creativity in recent creative text generation models such as those for automatic poetry generation, despite renewed interest in the context of recent LLMs franceschelli2023creativity. In fact, existing automatic poetry generation models are typically not evaluated regarding how different generated poems are from existing poems in the training set but with the Turing test: can humans distinguish whether a poem is human authored or automatically generated (hopkins-kiela-2017-automatically; lau2018deep; Manjavacas2019ARS)? However, this form of Turing test and other similar forms of human evaluation may contain an overlooked risk of failure: namely, if the automatically generated instances are (near-)copies of training data instances.
In this work, we fill this gap and evaluate, for the first time, automatic poetry generation systems for their diversity. As human evaluation is generally not well suited to assess diversity hashimoto-etal-2019-unifying, we automatically measure diversity by comparing distributions of generated and existing poems along formal, semantic and stylistic dimensions. This yields much better evidence of the models’ creative capabilities in contrast to being mere ‘stochastic parrots’.
Our main contributions are: (i) we conceptualize diversity of poetry generation systems along different dimensions: diversity on the structural, lexical, semantic and stylistic level; (ii) we assess different types of automatic poetry generation systems for diversity: general purpose word and character-level LLMs, both unconditioned and style-conditioned ones, on the one hand, and poetry-specific models, on the other hand; (iii) we evaluate each class of model for diversity across the different dimensions, by comparing the distribution of the human authored training data set to the distribution of generated poems. We find that on a distributional level, generated poems are considerably different from human ones. Concerning general purpose LLMs, some of them exhibit very high risk of memorization — an extreme form of lack of diversity — and this depends on the size of the training data set, the size and type of the LLM, and the type of training, as we show. Character-level style-conditioned general-purpose LLMs are most diverse.
Our work prepares the groundwork for truly creative generative AI models (veale2020leaps) and also has implications for the detection of generative AI (Sadasivan2023CanAT).
@@ 2 Related Work
Building systems able to generate diverse output has been a long-standing concern in NLG research (reiter-sripada-2002-squibs; van-deemter-etal-2005-squibs; foster-white-2007-avoiding) and remains a central issue in neural NLG (holtzman2019curious). The need for careful analysis of NLG systems diversity – beyond an assessment of the quality or fluency of single-best generation outputs – has been widely acknowledged (gatt2018survey; hashimoto-etal-2019-unifying; mahamood-zembrzuski-2019-hotel; celikyilmaz2020evaluation; tevet-berant-2021-evaluating; schuz-etal-2021-diversity). A well-known finding from this line of research is that neural NLG systems typically face a quality-diversity trade-off (ippolito-etal-2019-comparison; caccia2020; wiher2022decoding): their outputs are either well-formed and fluent or diverse and variable.
Work on evaluating diversity of NLG typically uses automatic metrics that quantify to what extent different outputs by the same system vary (hashimoto-etal-2019-unifying). In practice, though, evaluations of diversity in NLG differ widely across tasks (tevet-berant-2021-evaluating) and even adopt different notions of diversity (zarriess-etal-2021-decoding). At the same time, most of these notions focus on lexical or semantic aspects of diversity, e.g., local lexical diversity. For instance, ippolito-etal-2019-comparison compare decoding methods in dialog generation and image captioning, assessing lexical overlaps in -best NLG outputs for the same input. Global lexical diversity, on the other hand, measures whether the NLG system generates different outputs for different inputs. For instance, van-miltenburg-etal-2018-measuring define the global diversity of image captioning systems as their ability to generate different captions for a set of inputs, using metrics like the number of types in the output vocabulary, type-token ratio, and the percentage of novel descriptions. Similarly, hashimoto-etal-2019-unifying view diversity as related to the model’s ability to generalize beyond the training set, i.e., generate novel sentences.
Besides lexical diversity, work on open-ended or creative text generation tasks has been interested in diversity at a more general semantic level. For instance, zhang2018 and stasaski-hearst-2022-semantic aim at building dialogue systems that generate entertaining and semantically diverse responses in chit-chat dialog, where the goal is to avoid “safe and bland” responses that “average out” the sentences observed in the training set. Here, semantic diversity has been measured, e.g., with the help of embedding-based similarity (du-black-2019-boosting).
In our work on diversity in poetry generation, we complement these lexical and semantic aspects of diversity with aspects of formal diversity. We thus explore whether automatic poetry generation systems are able to capture the ‘full bandwidth’ of realizations of poetry found in the data distribution with which they have been trained, focusing mostly on global diversity.
@@
Automatic poetry generation is a long standing dream of AI research, dating back at least to the mid 20th century (e.g., Theo Lutz’ Stochastische Texte). While early modern systems were heavily hand-engineered gervas2001expert, more recent approaches are all trained on collections of human poetry lau2018deep; jhamtani2019learning; agarwal-kann-2020-acrostic but still extensively utilize human guidance e.g. to enforce formal characteristics of poetry such as rhyming wockener-etal-2021-end. belouadi2022bygpt5 have recently released a character-level decoder-only LLM (ByGPT5) capable of learning style-constraints such as rhyming without human involvement in model design.
@@ 3 Diversity In Poetry Generation
We first conceptualize diversity in poetry generation using formal and semantic criteria. As our dataset, we use QuaTrain (belouadi2022bygpt5) consisting of quatrains (in English and German). We describe it in more detail in §5.1 below.
@@
In poetry, as in other forms of art, creativity (sternberg1999handbook) plays a central role. A basic aspect of creativity is the models’ ability to generate poems that are different from the training data, i.e. have not been memorized as a whole. Therefore, we consider a low or minimal degree of memorization as a pre-requisite for diversity and analyze the portion of generated poems that are (near-)copies from the training data. To examine memorization, we proceed as in belouadi2022bygpt5. We apply the Ratcliff-Obershelp similarity (ratcliff1988pattern) to compare each poem in a sample with poems in the training corpus. If a generated quatrain exhibits a similarity score of 0.7 with a quatrain in the training data, we classify it as memorized. A quatrain can be divided into 4 verses or 2 couplets; thus, we also inspect memorization at the verse and couplet levels by comparing each verse or couplet in a sample to those in the training data. Higher thresholds for classification are used for these finer-grained comparison levels, as shorter texts have higher chances of being more similar in general. 11todo: 1YC: maybe add a table in the appendix to demonstrate the couplets/verses classified as memorized with threshold 0.7. Specifically, a verse with a similarity score 0.9 or a couplet 0.8 is considered as memorized. We define the memorization score of a sample as the proportion of memorized quatrains in that sample. How much LLMs memorize from their training data has been a question of central concern recently mccoy-etal-2023-much.
@@
Within a sample of generated poems, we consider differences at the level of poem length, i.e., their number of tokens, as a basic aspect of diversity at the formal or structural level. We analyze to what extent the length distribution of generated poems differs from the distribution in the training data. We define the length of a quatrain as the number of tokens contained: we eliminate all punctuation symbols and split the remaining text by white space. We report mean length, standard deviation, minimal and maximal length of samples. We additionally deploy distance measures between training data distribution and generated samples, in particular, a metric called histogram intersection (swain1991color), which measures the intersection area of two normalized histograms (and therefore returns values between 0 and 1).
@@
As a more complex dimension of formal diversity, we consider rhyming as a central aspect that characterizes the structure of a poem. Diversity can then be assessed by comparing rhyme distributions between generated samples and training data. In order to classify rhymes in our samples, we use the same classifier used to annotate QuaTrain. We distinguish between true rhymes, which involve different words, and repetitions, which refer to rhymes based on the same word.
@@
Lexical diversity is a standard aspect of diversity evaluation in NLG and is used to assess how generation outputs vary in their vocabulary, either at the local text level or at the global corpus level. We use the following metrics to measure the lexical diversity for both the training data and the generated samples: (i) Averaged type token ratio (ATTR). We calculate ATTR as the average of all type token ratios (richards1987type) (TTRs) for each quatrain in a sample, i.e. as a measure of local lexical diversity. (ii) Moving average type token ratio (MATTR). The MATTR (covington2010cutting) acts on the corpus level and calculates a moving average by sliding through the corpus using a window of fixed size. We deploy this metric as a measure of global lexical diversity. (iii) Measure of textual, lexical diversity (MTLD). The MTLD (mccarthy2005assessment) is calculated as the average length of a substring that maintains a specified TTR level. MTLD is deployed to measure lexical diversity on a global scale.
@@
Even if a poetry generation system does not directly copy data from the training data, the generated poems may still be semantically very similar to the training data distribution. We employ a multilingual distilled version of Sentence-BERT (SBERT) reimers-gurevych-2019-sentence as dense vector representations to measure semantic similarity between poems: (i) across the human train set and the generated poems, (ii) within human and generated poems. In particular, for each generated quatrain, we note down the similarity value of the most similar human quatrain, then report the average over all those maximum similarity values. We proceed analogously within the human training data and within the automatically generated poems.
4 Models
Our experiments use 2 different model classes.
4.1 Poetry-specific models
Deepspeare.
Deepspeare lau2018deep is specifically designed for poetry generation. Its core architecture consists of an LSTM language model, a pentameter model (specifically designed to learn iambic meter) and a rhyme model. During training, it takes sonnets as input data (three quatrains followed by a couplet) but ultimately processes the contained quatrains by splitting any given sonnet. The rhyme model processes ending words of quatrain verses and uses a margin-based loss to discriminate between rhyming and non-rhyming words. It is not limited to specific rhyme patterns but assumes that rhymes exist in the data. At inference time, Deepspeare generates quatrains.
Structured Adversary.
Like Deepspeare, Structured Adversary (SA) jhamtani2019learning incorporates different components: an LSTM language model and a discriminator used to decide whether line endings are typical for poetry. Both components are organized in an adversarial setup, where the language model acts as a generator, trying to generate poems that are misclassified by the discriminator, while the discriminator is trained to distinguish generated poems from real ones. SA is trained with sonnets as input data. At inference time, it generates quatrains.
4.2 General purpose LLMs
All models in this category are decoder-only transformer architectures. In our experiments, we train them in an unconditioned and style-conditioned manner (see Section 5.2).
GPT2
GPT2 radford2019language is the last GPT model made publicly available. It is a large word level transformer-based language model pretrained on approximately 40 GB of text. Four different model versions were released, with the number of parameters ranging from 125 million to 1.5 billion for the largest. In this work, we deploy two model versions: GPT2-small (125M parameters) and GPT2-large (774M parameters) for both English and German.
GPTneo
GPTneo (black-etal-2022-gpt) is an open-source token level LLM by EleutherAI (https://www.eleuther.ai/) with the aim to provide publicly available replications of GPT3. It is pre-trained on 825 GB of text data. Currently, four versions have been released, with the number of parameters ranging from 125 million up to 20 billion. We deploy GPTneo-small and GPTneo-xl with 125M and 1.3B parameters for English. GPTneo is not available for German.
LLaMA2/3
ByGPT5
ByGPT5 (belouadi2022bygpt5) is a decoder-only character level LLM based on the encoder-decoder character level model byT5 (xue-etal-2022-byt5) where the encoder part of byT5 is completely removed, reducing the number of parameters by 75%. The remaining decoder-only model is then pretrained using OpenWebText for English (38GB text data) and CC100 (conneau-etal-2020-unsupervised) (67GB text data) for German. Three versions are released for both English and German, with model sizes ranging from 73 to 298M parameters. We use ByGPT5-base (140M params) and ByGPT5-medium (290M) for both English and German.
5 Experimental Setup
5.1 Training Data
@@
We use QuaTrain, a large dataset of quatrains published by belouadi2022bygpt5. It consists of English and German quatrains and has been generated by aggregating different publicly available poetry datasets. QuaTrain contains human written quatrains but mixes them synthetically: every sequence of four consecutive lines from the underlying human data are included in order to increase dataset size. QuaTrain is automatically annotated for meter and rhyme using high-quality classifers (especially for rhyme). 22todo: 2YC: @Hannes, could you please add here how you preprocessed the data to drop the redundant ones? H: on it! Because our focus lies on the diversity of model outputs, we have to avoid repetitions in the training data created by the data augmentation methods used in its creation. To avoid lines appearing multiple times, we first parse the dataset sequentially, eliminating quatrains that overlap the preceding one. Because this method does not eliminate all overlaps, we then use an heuristic, deleting the ten percent of the quatrains which have the biggest overlap with other quatrains until there is no overlap remaining.
@@
We use a ratio of 9:1 to split the data into train and dev sets; we do not keep a test set since no held-out human data is needed for generation or evaluation. Further, as DeepSpeare and SA are designed to process sonnets and/or limerick data, we create pseudo sonnets for them using the corresponding data splits. Specifically, for each sonnet, we randomly draw three quatrains and one couplet from the QuaTrain datasets, ensuring that each comes from a different original quatrain. Table 1 provides the resulting data statistics.
5.2 Training
DeepSpeare.
DeepSpeare (lau2018deep) leverages pretrained static word vectors. We use QuaTrain and SonNet to train our own Word2vec embeddings (mikolov2013efficient) and the final sonnet models respectively. For the sonnet model training, we use a batch size of 128 and apply early stopping with a patience of 5 epochs; default settings are maintained for the other hyperparameters.
SA.
Unconditioned LLMs.
In this setup, we fine-tune our decoder-only LLMs in an unconditioned manner: we process quatrains during training without passing any information about rhyme (or meter). All models except GPTneo (being available only in English) are trained both in English and German. We fine-tune all models (English and German) for 50 epochs, with early stopping using a patience of 3 epochs.
Style-conditioned LLMs.
In contrast to unconditioned training, we provide information about rhyme (and meter) by prepending special style tokens to each quatrain during training. This follows the setup of belouadi2022bygpt5 and makes models explicitly aware of different rhyme schemes. As for the unconditioned variants, all models except GPTneo are trained in English and German. We use the same hyperparameter settings as for the unconditioned trained LLMs.
Summary.
We end up with 36 models that can be assigned to three categories: 1) Poetry specific LSTM-based models (DeepSpeare and SA). Besides a language model part, these models incorporate additional specialized components to handle poetry-specific stylistic features such as rhyme. We have 2 models for each language. 2) Unconditioned LLMs (transformer-based decoder-only general purpose LLMs). These models do not possess any specialized architecture for poetry. No information about meter or rhyme has actively been passed during training. We have two subcategories: word and character level models. The first group (the GPT and LLaMA series) processes data on the word/subword level. ByGPT5 represents the character-level group. We have 9 models for English and 7 for German. 3) Style-conditioned LLMs. These have the same architecture, models, and subgroups as category 2. Information about rhyme (and meter) is passed in the form of special tokens during training (only). We use the following notation to distinguish the size and training manners of the LLMs: a subscript of S/L indicates whether it is a smaller/larger version, and a superscript of “con” stands for conditioned training. E.g., GPT2 and GPT2 represent the unconditioned and conditioned trained GPT2 small models, respectively.
Table 4 (appendix) provides an overview of all models belonging to the second and third category (transformer-based LLMs).
5.3 Sampling

### v2

@@ Abstract
Natural Language Generation (NLG), and more generally generative AI, are among the currently most impactful research fields. Creative NLG, such as automatic poetry generation, is a fascinating niche in this area. While most previous research has focused on forms of the Turing test when evaluating automatic poetry generation — can humans distinguish between automatic and human generated poetry — we evaluate the diversity of automatically generated poetry (with a focus on quatrains), by comparing distributions of generated poetry to distributions of human poetry along structural, lexical, semantic and stylistic dimensions, assessing different model types (word vs. character-level, general purpose LLMs vs. poetry-specific models), including the very recent LLaMA3-8B, and types of fine-tuning (conditioned vs. unconditioned). We find that current automatic poetry systems are considerably underdiverse along multiple dimensions — they often do not rhyme sufficiently, are semantically too uniform and even do not match the length distribution of human poetry. Our experiments reveal, however, that style-conditioning and character-level modeling clearly increases diversity across virtually all dimensions we explore. Our identified limitations may serve as the basis for more genuinely diverse future poetry generation models.111Code + data: https://github.com/hgroener/diversity_in_poetry_generation
@@
Yanran Chen1, Hannes Gröner2, Sina Zarrieß2, Steffen Eger1 1 NLLG, University of Mannheim & University of Technology Nuremberg (UTN); https://nl2g.github.io/ 2 Computational Linguistics, Bielefeld University yanran.chen@uni-mannheim.de {hannes.groener,sina.zarriess}@uni-bielefeld.de steffen.eger@utn.de
@@ 1 Introduction
A key aspect of creative language generation is the ability to create new, original and interesting text, cf. Colton et al. (2012); Gatt and Krahmer (2018); Yi et al. (2020); Elgammal et al. (2017). To date, extremely little attention has been given to the evaluation of originality and creativity in recent creative text generation models such as those for automatic poetry generation, despite renewed interest in the context of recent LLMs Franceschelli and Musolesi (2023). In fact, existing automatic poetry generation models are typically not evaluated regarding how different generated poems are from existing poems in the training set but with the Turing test: can humans distinguish whether a poem is human authored or automatically generated (Hopkins and Kiela, 2017; Lau et al., 2018; Manjavacas et al., 2019)? However, this form of Turing test and other similar forms of human evaluation may contain an overlooked risk of failure: namely, if the automatically generated instances are (near-)copies of training data instances.
In this work, we fill this gap and evaluate, for the first time, (fine-tuned) automatic poetry generation systems in terms of their diversity. As human evaluation is generally not well suited to assess diversity Hashimoto et al. (2019), we automatically measure diversity by comparing distributions of generated and existing poems along formal, semantic and stylistic dimensions. This yields much better evidence of the models’ creative capabilities in contrast to being mere ‘stochastic parrots’.
Our main contributions are: (i) we conceptualize diversity of poetry generation systems along different dimensions: diversity on the structural (e.g., length), stylistic (e.g., rhyming), lexical and semantic level; (ii) we assess different types of automatic poetry generation systems for diversity: general purpose word- and character-level LLMs, both unconditioned and style-conditioned ones, on the one hand, and poetry-specific models, on the other hand; (iii) we evaluate each class of model for diversity across the different dimensions, by comparing the distribution of the human authored training data set to the distribution of generated poems. We find that on a distributional level, generated poems are considerably different from human ones. Character-level style-conditioned general-purpose LLMs are most diverse.
@@ 2 Related Work
Building systems able to generate diverse output has been a long-standing concern in NLG research (Reiter and Sripada, 2002; van Deemter et al., 2005; Foster and White, 2007) and remains a central issue in neural NLG (Holtzman et al., 2019). The need for careful analysis of NLG systems’ diversity – beyond an assessment of the quality or fluency of single-best generation outputs – has been widely acknowledged (Gatt and Krahmer, 2018; Hashimoto et al., 2019; Mahamood and Zembrzuski, 2019; Celikyilmaz et al., 2020; Tevet and Berant, 2021; Schüz et al., 2021). A well-known finding from this line of research is that neural NLG systems typically face a quality-diversity trade-off (Ippolito et al., 2019; Caccia et al., 2020; Wiher et al., 2022): their outputs are either well-formed and fluent or diverse and variable.
Work on evaluating diversity of NLG typically uses automatic metrics that quantify to what extent different outputs by the same system vary (Hashimoto et al., 2019). In practice, though, evaluations of diversity in NLG differ widely across tasks (Tevet and Berant, 2021) and even adopt different notions of diversity (Zarrieß et al., 2021). At the same time, most of these notions focus on lexical or semantic aspects of diversity, e.g., local lexical diversity. For instance, Ippolito et al. (2019) compare decoding methods in dialogue generation and image captioning, assessing lexical overlaps in -best NLG outputs for the same input. Chakrabarty et al. (2022) simply measure the local lexical diversity in automatic generated poems in terms of distinct unigrams. Global lexical diversity, on the other hand, measures whether the NLG system generates different outputs for different inputs. For instance, van Miltenburg et al. (2018) define the global diversity of image captioning systems as their ability to generate different captions for a set of inputs, using metrics like the number of types in the output vocabulary, type-token ratio, and the percentage of novel descriptions. Similarly, Hashimoto et al. (2019) view diversity as related to the model’s ability to generalize beyond the training set, i.e., generate novel sentences.
Besides lexical diversity, work on open-ended or creative text generation tasks has been interested in diversity at a more general semantic level. For instance, Zhang et al. (2018) and Stasaski and Hearst (2022) aim at building dialogue systems that generate entertaining and semantically diverse responses in chit-chat dialog. Here, semantic diversity has been measured, e.g., with the help of embedding-based similarity (Du and Black, 2019).
In our work on diversity in poetry generation, we complement both lexical and semantic aspects of diversity with aspects of formal diversity. We thus explore whether automatic poetry generation systems are able to capture the ‘full bandwidth’ of realizations of poetry found in the data distribution with which they have been trained, focusing mostly on global diversity.
@@
Automatic poetry generation is a long standing dream of AI research, dating back at least to the mid 20th century (e.g., Theo Lutz’ Stochastische Texte). While early modern systems were heavily hand-engineered Gervás (2001), more recent approaches are all trained on collections of human poetry Lau et al. (2018); Jhamtani et al. (2019); Agarwal and Kann (2020) but still extensively utilize human guidance e.g. to enforce formal characteristics of poetry such as rhyming Wöckener et al. (2021). Belouadi and Eger (2023) have recently released a character-level decoder-only LLM (ByGPT5) capable of learning style-constraints such as rhyming without human involvement in model design. Chakrabarty et al. (2022) propose a collaborative system for poetry, which can follow human instructions to write poems. They measure creativity of the generated poems via crowd workers, who decide which of two poems is more creative. While Chakrabarty et al. (2022) do not define creativity, it could be considered as generating novel poems outside the training data set; in contrast, we measure diversity by assessing whether poetry generation systems generate outputs that are as diverse as their human training data.
@@ 3 Diversity in Poetry Generation
We first conceptualize diversity in poetry generation using formal and semantic criteria.
@@
In poetry, as in other forms of art, creativity (Sternberg, 1999) plays a central role. A basic aspect of creativity is the models’ ability to generate poems that are different from the training data, i.e. have not been memorized as a whole. To examine memorization, we proceed as in Belouadi and Eger (2023). We apply the Ratcliff-Obershelp similarity (Ratcliff et al., 1988) to compare each poem in a sample with poems in the training corpus. If a generated quatrain exhibits a similarity score of 0.7 with a quatrain in the training data, we classify it as memorized. A quatrain can be divided into 4 verses or 2 couplets; thus, we also inspect memorization at the verse and couplet levels by comparing each verse or couplet in a sample to those in the training data. Higher thresholds for classification are used for these finer-grained comparison levels, as shorter texts have higher chances of being more similar in general. Specifically, a verse with a similarity score 0.9 or a couplet 0.8 is considered as memorized. We define the memorization score of a sample as the proportion of memorized quatrains in that sample. How much LLMs memorize from their training data has been a question of central concern recently McCoy et al. (2023).
@@
Within a sample of generated poems, we consider differences at the level of poem length, i.e., their number of tokens, as a basic aspect of diversity at the formal or structural level. We analyze to what extent the length distribution of generated poems differs from the distribution in the training data. We define the length of a quatrain as the number of tokens contained: we eliminate all punctuation symbols and split the remaining text by white space. We report mean length, standard deviation, minimal and maximal length of samples. We additionally deploy distance measures between training data distribution and generated samples, in particular, a metric called histogram intersection (Swain and Ballard, 1991), which measures the intersection area of two normalized histograms (and therefore returns values between 0 and 1).
@@
As a more complex dimension of formal diversity, we consider rhyming as a central aspect that characterizes the structure of a poem. Diversity can then be assessed by comparing rhyme distributions between generated samples and training data. In order to classify rhymes in our samples, we use the same classifier used to annotate QuaTrain (Belouadi and Eger, 2023). We distinguish between true rhymes, which involve different words, and repetitions, which refer to rhymes based on the same word.
@@
Lexical diversity is a standard aspect of diversity evaluation in NLG and is used to assess how generation outputs vary in their vocabulary, either at the local text level or at the global corpus level. We use the following metrics to measure the lexical diversity for both the training data and the generated samples: (i) Averaged type token ratio (ATTR). We calculate ATTR as the average of all type token ratios (Richards, 1987) (TTRs) for each quatrain in a sample, i.e. as a measure of local lexical diversity. (ii) Moving average type token ratio (MATTR). The MATTR (Covington and McFall, 2010) acts on the corpus level and calculates a moving average by sliding through the corpus using a window of fixed size. We deploy this metric as a measure of global lexical diversity. (iii) Measure of textual, lexical diversity (MTLD). The MTLD (McCarthy, 2005) is calculated as the average length of a substring that maintains a specified TTR level. MTLD is deployed to measure lexical diversity on a global scale.
@@
Even if a poetry generation system does not directly copy data from the training data, the generated poems may still be semantically very similar to the training data distribution. We employ a multilingual distilled version of Sentence-BERT (SBERT) Reimers and Gurevych (2019) as dense vector representations to measure semantic similarity between poems: (i) across the human train set and the generated poems, (ii) within human and generated poems. In particular, for each generated quatrain, we note down the similarity value of the most similar human quatrain, then report the average over all those maximum similarity values. We proceed analogously within the human training data and within the automatically generated poems.
4 Experiment Setup
Data
@@ 4 Experiment Setup
We use the QuaTrain dataset published by Belouadi and Eger (2023), which consists of English and German quatrains from different publicly available poetry datasets. The dataset contains human written quatrains but mixes them synthetically: every sequence of four consecutive lines from the underlying human data are included in order to increase dataset size. Besides, it is automatically annotated for meter and rhyme using high-quality classifers (especially for rhyme). Because our focus lies on the diversity of model outputs, we have to avoid repetitions in the training data created by the data augmentation methods used in its creation. To avoid lines appearing multiple times, we first parse the dataset sequentially, eliminating quatrains that overlap the preceding one. Because this method does not eliminate all overlaps, we then use a heuristic, deleting the ten percent of the quatrains which have the biggest overlap with other quatrains until there is no overlap remaining. We refer to the resulting dataset (again) as QuaTrain.
@@
QuaTrain is split into train and dev sets using a ratio of 9:1; we do not keep a test set since no held-out human data is needed for generation or evaluation. Further, as some models used in this work are designed to process sonnets and/or limerick data, we create pseudo sonnets for them, denoted as SonNet. Specifically, for each sonnet, we randomly draw three quatrains and one couplet from the corresponding data split of QuaTrain, ensuring that each comes from a different original quatrain. Table 1 provides the data sizes.
| Class | Model | Smaller | Larger | Lang |
| Poetry- specific | DeepSpeare | - | - | de/en |
| SA | - | - | de/en | |
| Unconditioned / Conditioned LLMs | ByGPT5 | 140m | 290m | de/en |
| GPT2 | 117m | 774m | de/en | |
| GPTNeo | 125m | 1.3b | en | |
| LLaMA2 | 7b | 13b | de/en | |
| LLaMA3 | 8b | de/en | ||
Models
We use 2 different model classes:
Poetry-specific Models: We select two models that integrate LSTM language models with additional components to generate quatrains with rhymes. DeepSpeare (Lau et al., 2018) utilizes a pentameter model to learn iambic meter and a rhyme model to distinguish between rhyming and non-rhyming words. Structured Adversary (SA) (Jhamtani et al., 2019) learns to rhyme in an adversarial setup, where a language model aims to generate poems misclassified by the discriminator, while a discriminator is trained to differentiate between generated and real poems. Both models can take sonnets as input during training and output quatrains during inference. For more detailed model descriptions, see Appendix A.1.
General Purpose LLMs: We consider several decoder-only transformer-based models, encompassing both (sub)word- and character-level models, as well as older and very recent models. We choose two model families from the GPT series, GPT2 (Radford et al., 2019) and GPTNeo (Black et al., 2022) (a replicated version of GPT3 by EleutherAI222https://www.eleuther.ai/), two from the LLaMA series, LLaMA2 (Touvron et al., 2023) and LLaMA3 (AI@Meta, 2024), and the character-level ByGPT5 (Belouadi and Eger, 2023). Except for LLaMA3, we consider one smaller and one larger variant within each model family based on model size. We train each model in both unconditioned and conditioned manners, with rhymes and meters exposed during training in the latter case. We encode styles with special tokens during training and allow the models to predict the styles autonomously during inference. For all LLMs, we employ consistent decoding strategies for generation: we use the default settings of the LLaMA2 chat models on Hugging Face333https://huggingface.co/spaces/huggingface-projects/llama-2-7b-chat but limit the number of newly generated tokens to 100 for the word-level models and 300 for the character-level ByGPT5 models.
We end up with a total of 36 models for German and English, categorized into three groups: 1) poetry specific LSTM-based models, 2) unconditioned LLMs, and 3) conditioned LLMs, as summarized in Table 2. SonNet is used for training 1), while QuaTrain is used for 2) and 3), separately for each language. We train all models using early stopping based on the perplexity/loss observed in the dev sets (details see Appendix A.2), as overfitting may negatively bias certain metrics like memorization rates. To distinguish between the different sizes and training manners of the LLMs, we use the following notation: a subscript of S/L indicates whether it is a smaller/larger version, and a superscript of “con” stands for conditioned training. E.g., GPT2 and GPT2 represent the unconditioned and conditioned trained GPT2 small models, respectively.
| Memorization () | Length () | Rhyme () | ||||||
| DE | EN | DE | EN | DE | EN | |||
| Couplet | Verse | Couplet | Verse | |||||
| Poetry-specific | 0.0000 | 0.006 | 0.0000 | 0.0046 | 0.752 | 0.745 | 0.992 | 0.825 |
| Character-level | 0.0000 | 0.010 | 0.0000 | 0.0087 | 0.815 | 0.813 | 0.893 | 0.895 |
| Word-level | 0.0476 | 0.048 | 0.0005 | 0.0309 | 0.686 | 0.700 | 1.057 | 0.852 |
| Unconditioned | 0.0003 | 0.045 | 0.0006 | 0.0324 | 0.686 | 0.681 | 1.107 | 0.937 |
| Conditioned | 0.0004 | 0.028 | 0.0002 | 0.0194 | 0.760 | 0.769 | 0.913 | 0.785 |
| Larger | 0.0005 | 0.037 | 0.0005 | 0.0290 | 0.713 | 0.705 | 1.111 | 0.861 |
| Smaller | 0.0003 | 0.039 | 0.0003 | 0.0237 | 0.726 | 0.756 | 0.931 | 0.890 |
| Lexical () | Semantic () | |||||||||
| DE | EN | DE | EN | |||||||
| ATTR | MATTR | MTLD | ATTR | MATTR | MTLD | Within | Across | Within | Across | |
| Poetry-specific | 0.928 | 0.895 | 162.8 | 0.890 | 0.863 | 126.0 | 0.577 | 0.669 | 0.509 | 0.601 |
| Character-level | 0.915 | 0.886 | 166.7 | 0.837 | 0.818 | 83.4 | 0.582 | 0.678 | 0.522 | 0.610 |
| Word-level | 0.922 | 0.874 | 114.7 | 0.871 | 0.835 | 82.7 | 0.629 | 0.693 | 0.587 | 0.634 |
| Unconditioned | 0.919 | 0.875 | 125.9 | 0.854 | 0.818 | 75.2 | 0.613 | 0.688 | 0.580 | 0.632 |
| Conditioned | 0.921 | 0.880 | 133.2 | 0.873 | 0.845 | 90.6 | 0.619 | 0.688 | 0.571 | 0.627 |
| Larger | 0.932 | 0.890 | 143.9 | 0.873 | 0.837 | 84.1 | 0.613 | 0.689 | 0.571 | 0.626 |
| Smaller | 0.902 | 0.861 | 115.6 | 0.839 | 0.814 | 74.3 | 0.623 | 0.688 | 0.577 | 0.631 |
5 Evaluation
We first report the results of diversity evaluation in §5.1, which is our main focus, followed by an examination of the relationship between diversity and overall quality through human evaluation in §5.2.
5.1 Diversity Evaluation